# Beyond Detection: Exploring Evidence-based Multi-Agent Debate for Misinformation Intervention and Persuasion

### v1

@@ Conclusion
We present ED2D, an evidence-based multi-agent debate (MAD) framework for misinformation detection and persuasive debunking. Extending prior approaches, ED2D introduces an evidence retrieval module, allowing agents to incorporate verifiable external facts into the argumentation. This design not only improves factual accuracy but also enhances interpretability and user-facing persuasiveness. Results demonstrates that ED2D consistently outperforms other baselines across three real-world datasets, including the latest Snopes25 benchmark. A controlled user study further demonstrates that ED2D-generated explanations are as persuasive as expert-written fact-checking. However, our findings also reveal a dual-use risk: when ED2D is incorrect, it may still persuade users in the wrong direction. Overall, ED2D offers a scalable and interpretable approach for proactive misinformation intervention. Future research will explore cost-efficient scaling, real-time deployment, and safeguards against deceptive or adversarial use.

### v2

@@ Conclusion
In this paper, we present ED2D, an evidence-based multi-agent debate (MAD) framework for misinformation detection and persuasive debunking. ED2D augments structured debate with an evidence retrieval module that grounds agent arguments in verifiable facts, improving both accuracy and the interpretability of explanations. Across three real-world datasets, ED2D consistently outperforms strong baselines. A controlled user study shows that its explanations are as persuasive as expert fact-checks. However, the results also reveal a dual-use risk: when ED2D is wrong, it can still shift user beliefs in misleading directions. Future work will focus on more efficient scaling, richer reasoning beyond binary labels, and safeguards against deceptive or adversarial use.